# Non-Prehensile Tool-Object Manipulation by Integrating LLM-Based Planning and Manoeuvrability-Driven Controls

### v5

@@ 3.3 Tool-Object Manipulation in Constrained Environments
To further evaluate the performance of the model in application scenarios, different shapes of walls are constructed as shown in Fig. 8(d)–(e). Two walls are designed with 90-degree and 65-degree for the inner-angles. Maneuvering a hook within a confined space presents greater challenges compared to using a stick. Additionally, a Y-shaped hook proves unsuitable for dragging objects in tight quarters. Therefore, in this experimental study, we opt for a hook tool with a right arm to navigate effectively within the constrained environment. Similar to the previous results, Fig. 8(d)–(e) also implements the task planner successfully to decompose the task and applies the stepping controller for object manipulation. The tool first aligns its $\mathbf{s}^{\text{tip}}$ to the first segment of the wall and adopts the proposed non-prehensile stepping motion controller stated in (2.6). The manipulandum is dragged out from the confined area by alternating between the action of ‘repositioning’ and ‘rotation-dragging’.
During the pulsing manipulation, the manipulandum maintains contact with the highest manoeuvrability region. The contact changes between the centre of the highest manoeuvrability region $\mathbf{p}_{\ast}$ with the manipulandum is visualized in Fig. 9(ii). The error between the $\mathbf{p}^{\text{obj}}$ and the wall exit for each experiment are minimised with time, as shown in Fig. 9.
3.4 Comparison
@@
We compare our system with other state-of-the-art methods. In terms of LLM-based task decomposition, we assess the success rates of our approach with zero-shot and few-shot learning methods [43], DELTA [21], SayPlan [22], and fine-tuning on a smaller dataset, as shown in Fig. 9(iv). In the comparison, zero-shot and few-shot learning refer to using prompts solely with a pre-trained model, rather than with a fine-tuned model. We consider task decomposition successful only if the output is optimal, with no unnecessary or redundant steps.
@@
We assess the tool analysis method by identifying the highest manoeuvrability point across 32 tool images, with the results outlined in Fig. 9(v) and 10. The center of the manipulated manipulandum is taken as the ground truth. For our analysis, we consider the average error, root mean square error (RMSE), and mean absolute error (MAE) as the key metrics. The results are visualized in Fig. 10, showcasing the differences between the ground truth and the computed results under various methodologies. In the comparison, we observe that the total variation regularization (TVR) method [44] had a relatively higher difference from the ground truth. The keypoint-inspired learning approach (similar to [45]) yields comparable results to our method. However, the keypoint approach requires manual labelling of large amounts of data and model training, and its accuracy is highly dependent on the quality of the dataset. As shown in Fig. 10, both the keypoint and our methods had lower errors along the x-axis than the y-axis. Overall, both achieved relatively lower errors than the TVR method. Yet, in general, our proposed method demonstrated more stable performance and higher accuracy in terms of manoeuvrability computation.
4 Conclusion
In this paper, we present a new manoeuvrability-driven approach for tool-object manipulation. The LLM is integrated for task decomposition, generating collaborative motion sequences for a dual-arm robot system. A compact geometrical-based affordance model for describing the potential functionality and computing the highest manoeuvrability region of a tool is developed. A non-prehensile motion controller and a stepping manipulation model are derived for TOM and incremental movements in a constrained area. Experimental results are reported and analysed for the proposed methodology validation. We illustrate the performance of the proposed methods in the accompanying video https://vimeo.com/917120431. Additional details of the LLMs and experiments are included in the supplementary materials of this paper.
Our method introduces a new affordance and manoeuvrability paradigm for tool-based object manipulation. To obtain a better performance, we split the model into task decomposition and mathematical motion models. However, the logical fault in the LLM’s response may be unseen and thus lead to inappropriate motion. In our experiments, there are a few times when the LLM presents infeasible plans. Moreover, the current affordance model presents promising results with simple geometrical shapes. Dynamic shapes like deformable objects may be complicated to perform accurate modelling. In terms of manoeuvrability, it may be complicated to compute an accurate result for scenes with unstable illumination, low contrast in images, large height differences in objects (tools and the manipulandum), etc. We simplified these cases using ArUco code for real-time object tracking in the experiments.
Appendix
4.1 LLM-Based High-Level Symbolic Task Planner
The training data for fine-tuning an LLM is generated with the pose of the robots, tools, block, and target randomly assigned based on the task description. The available motion functions are listed in Table 1.
Table 1: Motion Functions Available
Function Description
$\mathtt{grasp}(arm,tool)$ To grasp a tool with the robot arm.
$\mathtt{approach}(arm,tool,m)$ To approach the location of m with tool using arm.
$\mathtt{interact}(arm,tool,m,goal)$ To move m to the goal location with the tool.
$\mathtt{stepping}(arm,tool,m)$ To move m out from the bounded area with the tool of the arm through contact pulsing motions.
$\mathtt{pass}(arm1,tool,m,arm2)$ To pass m to another arm’s workspace.
$\mathtt{release}(arm,tool)$ To release the tool back to its original place with the arm.
Under various scenarios, the input and output of the task planner are illustrated in Fig. 11, 12, 13, where $x$ and $y$ refer to the coordinates of the items, such as the position of the block. The instructions and scene information are embedded to form the input for the planner. The scene information includes the positions of the robots, tools, block, target, and the key points of the tools and walls (if applicable).
Fig. 11: The input and output of the task planner for a long-horizon task with two tools. The input prompt is composed of the task instruction, the coordinates of robots, tools, and the block. The generated task sequence: first grasps both tools, then manipulates the block from right to the target location, and returns both tools to their original location.
Fig. 12: The input and output of the task planner for a long-horizon task with a single tool. The input prompt is composed of the task instruction, the coordinates of robots, tools, and the block. The generated task sequence: first grasps the tool with the right arm, then passes the block from the right to the left side, and puts the tool back to its original location once the block is at the target location.
Fig. 13: The input and output of the task planner with environmental constraints. The input prompt is composed of the task instruction, the coordinates of robots, tools, the block, and the wall. The generated task sequence: first grasps the hook, then incrementally controls the tool to drag the block out from the constrained area, and returns the hook to its original location.
4.2 Additional Experiments
Different Tools in a Long Horizon Task
  • 1.
    Case 1: Using a stick and then a hook to push the blue block to a faraway target, from right to left (see Fig. 14).
  • 2.
    Case 2: Using a stick and then a hook to push the blue block to a faraway top location, from left to top-right (see Fig. 15).
  • 3.
    Case 3: Using a Y-shaped hook and then a stick to push the blue block to the left-hand side target, from top-right to left (see Fig. 16).
Fig. 14: Case 1: (1) Observe the scene; (2) grasp and use the stick to push the blue block to the centre to pass it to the left arm; (3) use the hook to push the blue block to the target; (4) the blue block is at the desired location and both tools are put back to their original place.
Fig. 15: Case 2: (1) Observe the scene; (2) grasp and use the stick to push the blue block to the centre to pass it to the left arm; (3) use the top part of the hook to push the blue block to the top-right location; (4) the blue block is at the desired location and both tools are put back to their original place.
Fig. 16: Case 3: (1) Observe the scene; (2) the right arm grasps and uses the left side of the Y-shaped hook to push the blue block closer to the left arm; (3) the left arm uses the left side of the stick to continue pushing the blue block to the target location; (4) the blue block is at the desired location and the tools are returned to its original place.
Single Tool in a Long Horizon Task: Tool Sharing
  • 1.
    Case 4: Using a hook to push the blue block to a faraway target, from right to left (see Fig. 17).
  • 2.
    Case 5: Using a hook to push the blue block to the top-right location, from center-left to top-right (see Fig. 18).
Fig. 17: Case 4: (1) Observe the scene; (2) the right arm grasps and uses the hook to push the blue block to the centre to pass it to the left arm; (3) the right arm releases the hook and the left arm uses the inter part of the hook to manipulate the blue block to the target location; (4) the blue block is at the desired location and the hook is returned to its original place.
Fig. 18: Case 5: (1) Observe the scene; (2) the left arm grasps and uses the right side of the hook to push the blue block closer to the right arm; (3) the left arm releases the hook and the right arm uses the top part of the hook to push the blue block to the target location; (4) the blue block is at the desired location and the hook is returned to its original place.
Manipulation in a Constrained Environment
  • 1.
    Case 6: Using a hook to get the blue block out from a constrained environment (see Fig. 19).
  • 2.
    Case 7: Using a hook to get the blue block out from a constrained environment with a different configuration (see Fig. 20).
Fig. 19: Case 6: (1) Observe the scene; (2) the right arm grasps the hook and approaches the blue block by aligning the top part of the tool to the wall; (3) the blue block is dragging out slowly with the hook being repositioned and rotation-dragging; (4) the blue block out and the hook is returned to its original place.
Fig. 20: Case 7: (1) Observe the scene; (2) the right arm grasps the hook and approaches the blue block by aligning the top part of the tool to the wall; (3) the blue block is dragging out slowly with the hook being repositioned and rotation-dragging; (4) the blue block out and the hook is returned to its original place.
5 Comparative Analysis of Task Decomposition Methodologies
In this section, we present a comparative analysis of various methodologies employed for robot task planning, specifically focusing on their success rates in different task scenarios. The methodologies evaluated include Zero-shot learning, Few-shot learning, DELTA [21], SayPlan [22], planning domain definition language (PDDL) [46], behaviour tree [47], and Fine-tuning with 200 data, and our proposed approach. The results are summarized in Table 2. Scenarios Evaluated:
  • 1.
    Our experiment setting: The position of the robot, tools, block, and target are based on our experiment settings
  • 2.
    New Case 1: New language instruction with the positions are based on a slightly larger table and robot’s workspace settings.
  • 3.
    New Case 2: New language instruction with the positions are based on a random-sized table and the robot’s workspace settings.
@@ 5 Comparative Analysis of Task Decomposition Methodologies
Methods Our experiment setting New Case 1 New Case 2 Overall(%)
Zero-shot learning 0.05 - - 1.67%
Few-shot learning 0.30 0.24 0.13 22.3%
@@
Behaviour tree [47] 0.99 0.42 0.04 48.3%
Fine-tuning (200 data) 0.83 0.77 0.69 76.3%
@@
In our experiment setting, most methods performed well, with PDDL and Behavior Trees achieving the highest initial success rates. However, they struggled in more complex scenarios, indicating limited generalizability. In New Case 1, our proposed method maintained a high success rate of 0.97, while PDDL and Behavior Tree approaches dropped to 0.42, highlighting their challenges in adapting to new instructions. Specifically, when the setting was expanded to twice the original size, the rules and conditions in PDDL and Behavior Trees proved ineffective. For instance, one condition stipulated that if the distance between the block and the right arm was less than a specified amount, then the right arm would move. However, as the workspace was enlarged, these conditions remained based on the original setting, resulting in no response from the system in some situations since none of the conditions were fulfilled.
A similar trend was observed in New Case 2, where our approach achieved an impressive success rate of 0.95, while other methods demonstrated lower effectiveness and adaptability. In general, fine-tuning can effectively adapt to new settings primarily because it can interpret new instructions from the prompt input and adjust its output accordingly. Overall, the fine-tuning method achieves superior results due to its high adaptability and generalizability to new instructions.

### v6

@@ 3.3 Tool-Object Manipulation in Constrained Environments
To further evaluate the performance of the model in application scenarios, different shapes of walls are constructed as shown in Fig. 8(d)–(e). Two walls are designed with 90-degree and 65-degree for the inner-angles. Maneuvering a hook within a confined space presents greater challenges compared to using a stick. Additionally, a Y-shaped hook proves unsuitable for dragging objects in tight quarters. Therefore, in this experimental study, we opt for a hook tool with a right arm to navigate effectively within the constrained environment. Similar to the previous results, Fig. 8(d)–(e) also implements the task planner successfully to decompose the task and applies the stepping controller for object manipulation. The tool first aligns its $\mathbf{s}^{\text{tip}}$ to the first segment of the wall and adopts the proposed non-prehensile stepping motion controller stated in Sec. 2.6. The manipulandum is dragged out from the confined area by alternating between the action of ‘repositioning’ and ‘rotation-dragging’.
During the pulsing manipulation, the manipulandum maintains contact with the highest manoeuvrability region. The contact changes between the centre of the highest manoeuvrability region $\mathbf{p}_{\ast}$ with the manipulandum are visualized in Fig. 9(ii). The error between the $\mathbf{p}^{\text{obj}}$ and the wall exit for each experiment is minimised with time, as shown in Fig. 9.
3.4 Comparison and Analysis
@@
We compare our system with other state-of-the-art methods. In terms of LLM-based task decomposition, we assess the success rates of our approach with zero-shot and few-shot learning methods [4], DELTA [19], SayPlan [29], and fine-tuning on a smaller dataset, as shown in Fig. 10. In the comparison, zero-shot and few-shot learning refer to using prompts solely with a pre-trained model, rather than with a fine-tuned model. We consider task decomposition successful only if the output is optimal, with no unnecessary or redundant steps.
@@
Methods Our settings NC1 NC2 Overall(%)
Zero-shot 0.05 - - 1.67%
Few-shot 0.30 0.24 0.13 22.3%
@@
B. tree [2] 0.99 0.42 0.04 48.3%
FT (200) 0.83 0.77 0.69 76.3%
@@
To evaluate the generalizability of our framework against state-of-the-art methods, we conducted a comparative analysis of various approaches to robot task planning, focusing on their success rates in previously unseen task scenarios. The evaluated methods include Zero-shot learning, Few-shot learning, DELTA [19], SayPlan [29], planning domain definition language (PDDL) [12], behaviour tree [2], and Fine-tuning with 200 data (FT 200), and our proposed approach. The results are summarized in Table 2. Scenarios Evaluated: Our experiment settings: The position of the robot, tools, block, and target are based on our experiment settings; new case 1 (NC1): New language instruction with the positions are based on a slightly larger table and robot’s workspace settings; new case 2 (NC2): New language instruction with the positions are based on a random-sized table and the robot’s workspace settings.
For a fair comparison of generalization capability, all baseline methods, including PDDL and Behavior Tree, were evaluated under a fixed configuration across all scenarios. Specifically, no manual retuning or reconfiguration (e.g., workspace bounds, distance thresholds, condition triggers, or collision margins) was performed for NC1 and NC2. Therefore, NC1 and NC2 are zero-shot transfer settings for these methods, where the task instructions and workspace scale change without updating the underlying symbolic rules or conditions.
In our experiment setting, most methods performed well, with PDDL and Behavior Trees achieving the highest initial success rates. However, their performance degraded significantly in NC1 and NC2, indicating limited generalizability under zero-shot transfer. Specifically, when the workspace was expanded or resized, the symbolic rules and absolute geometric thresholds used in PDDL and Behavior Trees, which were designed for the original workspace, were no longer applicable. For example, one condition stipulated that if the distance between the block and the right arm was below a fixed threshold, the right arm would initiate motion. Once the workspace scale changed, such conditions were no longer satisfied, resulting in no valid actions being triggered in certain situations.
A similar trend was observed in NC2, where our approach achieved a success rate of 0.95, while other methods showed substantially reduced performance. In contrast, fine-tuned learning-based methods maintain higher success rates because they can interpret task instructions and scene layouts directly from the prompt input and adjust their outputs accordingly. Overall, our proposed method demonstrates higher adaptability and generalization to unseen task instructions and workspace variations.
Table 3: Error in Tool Maneuverability Points
Methods Average(%) RMSE(%) MAE(%)
TVR 38% 49% 23%
Keypoint 16% 18% 10%
Ours 13% 15% 8%
Fig. 12: Component-wise ablation on point selection with a fixed subtask sequence. Top row: (a) predefined fixed point with the proposed controller; (b) endpoint-based selection; (c) side-picker strategy; (d) geometric-center selection. Bottom row: evolution of the positional error between the object and the target for (a)–(d).
Fig. 13: Results obtained using straight-line and basic position control in constrained environment settings. Structural deformation of the tools and walls is observed. The red dashed lines denote the reference geometry from the leftmost frame, illustrating the displacement of the wall relative to its initial configuration.
We assess the tool analysis method by identifying the highest manoeuvrability point across 32 tool images, with the results outlined in Table 3 and Fig. 11. The center of the manipulated manipulandum is taken as the ground truth. For our analysis, we consider the average error, root mean square error (RMSE), and mean absolute error (MAE) as the key metrics. The results are visualized in Fig. 11, showcasing the differences between the ground truth and the computed results under various methodologies. In the comparison, we observe that the total variation regularization (TVR) method [26] had a relatively higher difference from the ground truth. The keypoint-inspired learning approach (similar to [22]) yields comparable results to our method. However, the keypoint approach requires manual labelling of large amounts of data and model training, and its accuracy is highly dependent on the quality of the dataset. As shown in Fig. 11, both the keypoint and our methods had lower errors along the x-axis than the y-axis. Overall, both achieved relatively lower errors than the TVR method. Yet, in general, our proposed method demonstrated more stable performance and higher accuracy in terms of manoeuvrability computation.
The proposed framework integrates LLM-based task decomposition, manoeuvrability-driven point selection, and a non-prehensile motion controller. To disentangle the contribution of individual components, we conduct a targeted ablation study by bypassing the LLM-based task decomposition and fixing the same subtask sequence across all trials. This allows us to isolate the effects of point selection and motion control on tool-object manipulation performance.
As shown in Fig.12(a), we first replace the manoeuvrability-based point selection with a predefined fixed contact point while retaining the proposed motion controller. We then evaluate several baseline point-selection strategies: (b) selecting a point at a fixed distance from the tool endpoint, (c) a side-picker strategy that only enforces the object to remain on the correct side of the tool, and (d) selecting the geometric center of the tool.
The manipulation outcomes and the corresponding evolution of the position error between the object and the target are visualized in Fig.12. With a fixed point, the controller can still manipulate the object stably and guide it toward the target. In contrast, the endpoint-based and side-picker strategies result in significant object slipping along the tool, leading to large final errors, as shown in Fig.12(b,c). The geometric-center strategy yields behavior closer to the fixed-point baseline but still exhibits noticeable drift from the tool center. These results indicate that effective point selection and a manoeuvrability-based controller are critical for stable non-prehensile manipulation.
We additionally evaluated simpler motion-control baselines in constrained environments, including straight-line pushing and basic position control. In these settings, such controllers frequently failed to maintain stable contact and resulted in collisions with surrounding boundaries, leading to task failure and potential damage to the environment and to the tool (see Fig. 13. These failures are primarily due to the strong reliance of simple controllers on accurately modeled workspace geometry and precise contact conditions. In contrast, the proposed stepping and rotation-dragging controller adapts its motion incrementally based on visual feedback, enabling safer and more robust interaction under spatial constraints without requiring an explicit or highly accurate environment model.
4 Discussion and Conclusion
In this paper, we present a new manoeuvrability-driven approach for tool-object manipulation. The LLM is integrated for task decomposition, generating collaborative motion sequences for a dual-arm robot system. A compact geometrical-based affordance model for describing the potential functionality and computing the highest manoeuvrability region of a tool is developed. A non-prehensile motion controller is developed for TOM, utilizing a stepping controller for incremental manipulation within a constrained environment. Experimental results are reported and analysed for the proposed methodology validation. We illustrate the performance of the proposed methods in the accompanying video https://vimeo.com/917120431.
4.1 Discussion
Our method introduces a new affordance and manoeuvrability paradigm for tool-based object manipulation. To improve performance, the framework is separated into task decomposition and analytical motion models. This modular design allows the LLM to handle high-level planning using cloud computation, while the local computer executes physically grounded analytical models for low-level motion. In addition, the non-prehensile stepping controller enables incremental manipulation of objects in constrained environments. Unlike approaches that require computing an optimal trajectory for dragging the object out of a confined space, this method performs iterative small adjustments, alternating between tool repositioning and incremental rotation-dragging of the object until it is fully extracted. This strategy allows stable and precise manipulation in highly restricted areas, inspired by animal tool-use behaviors, and supports real-time execution without the need for high-end local GPU resources.
However, the method has limitations. The LLM may occasionally generate infeasible plans, which can lead to inappropriate motions. To improve generalization and enhance transferability to unseen scenarios, future work will explore alternative strategies such as domain adaptation and transfer learning. Moreover, the current affordance model presents promising results with simple geometrical shapes. Dynamic shapes like deformable objects may be complicated to perform accurate modelling. Manoeuvrability computation can also be affected by unstable illumination, low contrast, or large height differences between objects. In our experiments, these challenges were mitigated using ArUco markers for real-time tracking.
4.2 Conclusion
This work presents a manoeuvrability-driven framework for tool-object manipulation that integrates LLM-based task decomposition, a geometrical affordance model, and a non-prehensile stepping controller for incremental manipulation in constrained environments. Experimental validation demonstrates the effectiveness of this approach for collaborative dual-arm tasks and various tool configurations.